# SAND: One-Shot Feature Selection with Additive Noise Distortion

### v1

@@ Abstract
where $x_{i}$ is the input feature, $\tilde{x}_{i}$ the output, $z_{i}$ a Gaussian noise, and $a_{i}$ trainable gain such that $\sum_{i}{a_{i}^{2}}=k$. This formulation induces an automatic clustering effect, driving $k$ of the $a_{i}$ gains to $1$ (selecting informative features) and the rest to 00 (discarding redundant ones) via weighted noise distortion and gain normalization. Despite its extreme simplicity, our method delivers state-of-the-art performance on standard benchmark datasets and a novel real-world dataset, outperforming or matching existing approaches without requiring hyperparameter search for $k$ or retraining. Theoretical analysis in the context of linear regression further validates its efficacy. Our work demonstrates that simplicity and performance are not mutually exclusive, offering a powerful yet straightforward tool for feature selection in machine learning.
@@ 2 Selection with Additive Noise Distortion (SAND)
We then feed $\tilde{\boldsymbol{\underline{x}}}$ to the neural network during the training phase instead of $\boldsymbol{\underline{x}}$ as illustrated in Figure 1(b)). Here, $\boldsymbol{\underline{z}}$ is a Gaussian vector with i.i.d. entries with zero-mean and standard deviation $\sigma$. In this setting, when $a_{i}$ is close to $1$, $\tilde{\boldsymbol{x}}_{i}$ is close to noiseless $\boldsymbol{x}_{i}$, and when $a_{i}$ is close to 00, $\tilde{\boldsymbol{x}}_{i}$ becomes almost pure noise (the signal-to-noise ratio is proportional to $\frac{a_{i}^{2}}{(1-a_{i})^{2}\sigma^{2}}$). During the training phase, we allow the $a_{i}$’s to be trained alongside the other parameters of the network. The architecture of the neural network and the loss function remain unchanged; the only difference is the addition of $n$ extra parameters ($a_{i}$’s) to optimize. As training progresses, we observe that $k$ of the $a_{i}$’s cluster around 1, indicating the selected features, while the remaining $a_{i}$’s cluster around 0, indicating the neglected features. We refer to this approach as SAND, which stands for Selection with Additive Noise Distortion.
@@
The first non-trivial property is that there is always an optimal $\underline{a}$ whose entries are between 00 and $1$. Here, optimal means with respect to the loss function of the network. To prove that, assume there is an optimal $\underline{a}$ that has an $a_{j}<0$. By replacing $a_{j}$ by $-a_{j}$, while the constraint (2) still holds, $\tilde{\boldsymbol{x}}_{j}$ is a less noisy version of $\boldsymbol{x}_{j}$. On the other hand, if there is an optimal $\underline{a}$ that has an $a_{j}>1$, we can decrease $a_{j}$ to $1$, which results in $\tilde{\boldsymbol{x}}_{j}$ becoming a noiseless copy of $\boldsymbol{x}_{j}$, and increase other $a_{i}$’s that are less than $1$ towards $1$ to satisfy the condition (2); it decreases the noise added to those features as well. Therefore, we can confine the search space of $\underline{a}$ to the vectors that have all entries between 00 and $1$, i.e., $\underline{a}=\left(a_{1},a_{2},\dots,a_{n}\right)^{\top}$ that have
@@
Now, we analyze what is happening during the training. Intuitively, a more informative feature $x_{j}$ will get a gain $a_{j}$ closer to $1$ so that it will be passed to the neural network with less noise. Due to the constraint (2), this automatically yields smaller gains $a_{j^{\prime}}$ for other features which are less informative. Consequently, the less informative features become noisier, which makes them even less informative and pushes them to get even smaller gains. This leaves more room, due to (2), for the more informative features to get their gains closer to $1$ and become less noisy. This reinforcing loop, summarized in Figure 2, results in polarization of the gains around $1$ and 00. Ideally, we will end up with a vector $\underline{a}$ which has $k$ entries equal to $1$, indicating the selected features, and the rest equal to 00, indicating the neglected features. However, since in practice the smallest values have not necessarily converged to absolute zero, at the end of the training phase, we keep the top $k$ gains intact and manually set the $n-k$ smallest gains to 00, effectively removing those features. Notably, features with small gains are noisy and hence inherently ignored by other parts of the neural network.
@@
Remember that we can confine the search space to the $a_{i}$’s between 00 and $1$. Hence, according to (12), we have $\lambda\geq 0$. Therefore, the term at the end of (14) achieves its minima when there are $k$ of $a_{i}$’s equal to $1$ and $n-k$ of them equal to 00, which completes the proof.
@@
  • $\sigma$ indicates how firmly we would like to restrict the number of features to $k$. A higher value of $sigma$ places greater emphasis on precisely achieving $k$ features, resulting in faster binarization (polarization toward 00 and $1$) of the gains ($a_{i}$’s).
@@
We will see in the experiments that the method is not sensitive to the choice of $\sigma$ and $\alpha$. In fact, setting $\sigma$ within the range of standard deviation of the input features, and $\alpha=2$, yields nearly optimal results across all datasets. Thus, there is no need to fine-tune $\sigma$ and $\alpha$.
@@ 3 Experiments
Figure 3: Test metrics on 9 datasets over 10 trials. The metric is accuracy ()($\uparrow$)( ↑ ) for all except MAE ()($\downarrow$)( ↓ ) for CA Housing being a regression problem. SA = Sequential Attention, LLY = Batch-wise Attenuation, GL = Group LASSO, and SL = Sequential LASSO.
We implemented a neural network with one hidden layer and a ReLU activation. Given the variation in hidden layer widths across cited works, we opted for a width equal to $n/3$, where $n$ represents the dimensionality of the input data. Please refer to Table 3 of Appendix A for a comprehensive overview of the nine datasets, the corresponding number of epochs and batch size used for training, along with the mean accuracy/absolute-error of the model with all features. [^2] The code to reproduce our experiments is available at https://anonymous.4open.science/r/SAND-6BB1
Our evaluation included a comparison between SAND and four established feature selection algorithms, namely Sequential Attention (Yasuda et al., 2023), Sequential LASSO (Luo & Chen, 2014), LLY (Liao et al., 2021), and Group LASSO (Zhao et al., 2015). [^3] These algorithms represent the top-performing methods reported in the literature. A comparison with other approaches (Atashgahi et al., 2022; Lemhadri et al., 2021; Yamada et al., 2020; Sokar et al., 2022) demonstrates that SAND achieves better performance. For all methods except ours, training comprised a feature selection phase followed by a fitting phase wherein the neural network was retrained from scratch on the selected features. As for SAND, the fitting phase was omitted and the weights learned during the selection phase were directly utilized for inference. In other words, the gains corresponding to the non-selected features where set to zero while keeping all other weights of the model intact. Hence, from this point of view, our method offers two key benefits. Firstly, it demands fewer epochs ($33\%$ fewer epochs in our experiments). Secondly, it provides a streamlined pipeline where both selection and inference are handled by the same model.
@@
To provide additional insights, we varied the number of selected features $k$ under consistent settings and evaluated performance on the test set. Results are shown in Figure 4. Notably, SAND demonstrates its strength in feature selection, showcasing results that outperform or are comparable to other methods across different feature counts. This advantage is particularly significant given the fact that the best method is changing from dataset to dataset. Thus, there is a need for algorithms that deliver value beyond marginal accuracy improvements, prioritizing enhancements in computational demand and simplicity—qualities that our method exemplifies. It is important to note that all experiments were conducted using fixed values of $\sigma=1.5$ and $\alpha=2$, without any fine-tuning.
Figure 4: Test metrics for different $k$’s. Accuracy ()($\uparrow$)( ↑ ) for all except MAE ()($\downarrow$)( ↓ ) for CA Housing.
@@
As discussed in Remark 2.2, the parameter $\sigma$ influences the rate of gain polarization. To demonstrate this, we trained SAND model on MICE dataset for 2000 epochs, using $\sigma$ values of 1.0 and 2.0, while keeping other settings fixed. We recorded the gains every 10 epochs. Figure 5 presents the sorted gains for selected epochs.We observe that while the gains tend to cluster around $1$ and 00 in both plots, this clustering occurs at a higher rate for a larger $\sigma$. Moreover, to assess SAND’s sensitivity to $\sigma$, we replicated the initial experiment while varying $\sigma\in\{1.0,1.5,2.0,2.5,3.0\}$. Results are presented in Table 1. As evident in the table, our approach demonstrates high robustness to the selection of $\sigma$, which underscores a positive aspect of the proposed method.
@@ 4 Summary and Future Works
  • On par with the state-of-the-are performance: Through extensive experiments, we showed that the proposed method has effectively state-of-the-art performance.
@@ Appendix A Experimental set-up
Table 3: Dataset characteristics, experiment parameters, and all-features performance metrics. [^5] The metric is classification accuracy ()($\uparrow$)( ↑ ) for all except MAE ()($\downarrow$)( ↓ ) for CA Housing. Our study utilizes the entire MNIST and Fashion datasets, unlike related works. Additionally, the Activity dataset sourced from (Lemhadri et al., 2021)’s Google Drive and (Yasuda et al., 2023)’s repository contains 10,299 samples, as opposed to the 5,744 samples reported in the referenced papers.
@@
Moreover, the experiments were executed on a machine equipped with an NVIDIA GeForce RTX 4090 GPU with 24GB of RAM, paired with an AMD Ryzen 9 5900X 12-Core Processor featuring 24 threads. The code to reproduce our experiments is available at https://anonymous.4open.science/r/SAND-6BB1.
@@ Appendix B Experimental results
Table 4: Test metrics over 10 trials (mean $\pm$ standard deviation): Accuracy ()($\uparrow$)( ↑ ) for all except MAE ()($\downarrow$)( ↓ ) for CA Housing.
Dataset SA LLY GL SL SAND
Mice Protein $0.986\pm 0.005$ $0.983\pm 0.011$ $0.988\pm 0.007$ $0.987\pm 0.007$ $\mathbf{0.988}\pm 0.006$
MNIST $0.961\pm 0.002$ $0.949\pm 0.003$ $0.930\pm 0.006$ $\mathbf{0.963}\pm 0.002$ $0.962\pm 0.002$
MNIST-Fashion $0.836\pm 0.003$ $0.830\pm 0.004$ $0.828\pm 0.004$ $0.829\pm 0.003$ $\mathbf{0.836}\pm 0.005$
ISOLET $\mathbf{0.927}\pm 0.004$ $0.908\pm 0.011$ $0.924\pm 0.003$ $0.924\pm 0.007$ $0.926\pm 0.005$
COIL-20 $0.996\pm 0.005$ $0.998\pm 0.003$ $0.998\pm 0.003$ $0.996\pm 0.003$ $\mathbf{0.998}\pm 0.003$
Activity $0.920\pm 0.009$ $0.894\pm 0.050$ $\mathbf{0.930}\pm 0.004$ $0.915\pm 0.005$ $0.923\pm 0.004$
Madelon $0.786\pm 0.008$ $\mathbf{0.862}\pm 0.027$ $0.612\pm 0.010$ $0.603\pm 0.011$ $0.732\pm 0.021$
HAR70 $0.901\pm 0.001$ $\mathbf{0.910}\pm 0.001$ $0.910\pm 0.003$ $0.908\pm 0.004$ $0.897\pm 0.005$
CA Housing $0.639\pm 0.063$ $0.636\pm 0.063$ $0.590\pm 0.097$ $0.589\pm 0.097$ $\mathbf{0.521}\pm 0.027$
@@
Table 5: Selection consistency on HAR70: feature selection frequency of the useful features $(0-5)$ and the misselected ones over 10 runs for $k=6$.
Feature SA SL SAND GL LLY
0 10 10 10 10 10
1 8 9 8 9 10
2 10 10 10 10 10
3 10 10 10 10 10
4 10 7 10 10 10
5 2 10 10 10 10
Misselected 10 4 2 1 0

### v2

@@ Abstract
where $x_{i}$ is the input feature, $\tilde{x}_{i}$ the output, $z_{i}$ a Gaussian noise, and $a_{i}$ trainable gain such that $\sum_{i}{a_{i}^{2}}=k$. This formulation induces an automatic clustering effect, driving $k$ of the $a_{i}$ gains to $1$ (selecting informative features) and the rest to 0 (discarding redundant ones) via weighted noise distortion and gain normalization. Despite its extreme simplicity, our method achieves competitive performance on standard benchmark datasets and a novel real-world dataset, often matching or exceeding existing approaches without requiring hyperparameter search for $k$ or retraining. Theoretical analysis in the context of linear regression further validates its efficacy. Our work demonstrates that simplicity and performance are not mutually exclusive, offering a powerful yet straightforward tool for feature selection in machine learning.
@@ 2 Selection with Additive Noise Distortion (SAND)
We then feed $\tilde{\boldsymbol{\underline{x}}}$ to the neural network during the training phase instead of $\boldsymbol{\underline{x}}$ as illustrated in Figure 1(b)). Here, $\boldsymbol{\underline{z}}$ is a Gaussian vector with i.i.d. entries with zero-mean and standard deviation $\sigma$. In this setting, when $a_{i}$ is close to $1$, $\tilde{\boldsymbol{x}}_{i}$ is close to noiseless $\boldsymbol{x}_{i}$, and when $a_{i}$ is close to 0, $\tilde{\boldsymbol{x}}_{i}$ becomes almost pure noise (the signal-to-noise ratio is proportional to $\frac{a_{i}^{2}}{(1-a_{i})^{2}\sigma^{2}}$). During the training phase, we allow the $a_{i}$’s to be trained alongside the other parameters of the network. The architecture of the neural network and the loss function remain unchanged; the only difference is the addition of $n$ extra parameters ($a_{i}$’s) to optimize. As training progresses, we observe that $k$ of the $a_{i}$’s cluster around 1, indicating the selected features, while the remaining $a_{i}$’s cluster around 0, indicating the neglected features. We refer to this approach as SAND, which stands for Selection with Additive Noise Distortion.
@@
The first non-trivial property is that there is always an optimal $\underline{a}$ whose entries are between 0 and $1$. Here, optimal means with respect to the loss function of the network. To prove that, assume there is an optimal $\underline{a}$ that has an $a_{j}<0$. By replacing $a_{j}$ by $a_{j}$, while the constraint (2) still holds, $\tilde{\boldsymbol{x}}_{j}$ is a less noisy version of $\boldsymbol{x}_{j}$. On the other hand, if there is an optimal $\underline{a}$ that has an $a_{j}>1$, we can decrease $a_{j}$ to $1$, which results in $\tilde{\boldsymbol{x}}_{j}$ becoming a noiseless copy of $\boldsymbol{x}_{j}$, and increase other $a_{i}$’s that are less than $1$ towards $1$ to satisfy the condition (2); it decreases the noise added to those features as well. Therefore, we can confine the search space of $\underline{a}$ to the vectors that have all entries between 0 and $1$, i.e., $\underline{a}=\left(a_{1},a_{2},\dots,a_{n}\right)^{\top}$ that have
@@
Now, we analyze what is happening during the training. Intuitively, a more informative feature $x_{j}$ will get a gain $a_{j}$ closer to $1$ so that it will be passed to the neural network with less noise. Due to the constraint (2), this automatically yields smaller gains $a_{j^{\prime}}$ for other features which are less informative. Consequently, the less informative features become noisier, which makes them even less informative and pushes them to get even smaller gains. This leaves more room, due to (2), for the more informative features to get their gains closer to $1$ and become less noisy. This reinforcing loop, summarized in Figure 2, results in polarization of the gains around $1$ and 0. Ideally, we will end up with a vector $\underline{a}$ which has $k$ entries equal to $1$, indicating the selected features, and the rest equal to 0, indicating the neglected features. However, since in practice the smallest values have not necessarily converged to absolute zero, at the end of the training phase, we keep the top $k$ gains intact and manually set the nkn-k smallest gains to 0, effectively removing those features. Notably, features with small gains are noisy and hence inherently ignored by other parts of the neural network.
@@
Remember that we can confine the search space to the $a_{i}$’s between 0 and $1$. Hence, according to (12), we have $\lambda\geq 0$. Therefore, the term at the end of (14) achieves its minima when there are $k$ of $a_{i}$’s equal to $1$ and nkn-k of them equal to 0, which completes the proof.
@@
  • $\sigma$ indicates how firmly we would like to restrict the number of features to $k$. A higher value of sigmasigma places greater emphasis on precisely achieving $k$ features, resulting in faster binarization (polarization toward 0 and $1$) of the gains ($a_{i}$’s).
@@
We will see in the experiments that the method is not sensitive to the choice of $\sigma$ and $\alpha$. In fact, setting $\sigma$ within the range of standard deviation of the input features, and $\alpha=2$, yields nearly optimal results across all datasets. Thus, there is no need to fine-tune $\sigma$ and $\alpha$. Moreover, to ensure stable training, we apply clipping to the $a_{i}$ values throughout the training process, keeping them within the range $[0,1]$.
@@ 3 Experiments
Figure 3: Test metrics on 9 datasets over 10 trials. The metric is accuracy $(\uparrow)$ for all except MAE $(\downarrow)$ for CA Housing being a regression problem. SA = Sequential Attention, LLY = Batch-wise Attenuation, GL = Group LASSO, SL = Sequential LASSO, and STG = Stochastic Gates.
We implemented a neural network with one hidden layer and a ReLU activation. Given the variation in hidden layer widths across cited works, we opted for a width equal to $n/3$, where $n$ represents the dimensionality of the input data. Please refer to Table 3 of Appendix A for a comprehensive overview of the nine datasets, the corresponding number of epochs and batch size used for training, along with the mean accuracy/absolute-error of the model with all features. [^2] The code to reproduce our experiments is available at https://github.com/csem/SAND
Our evaluation included a comparison between SAND and five established feature selection algorithms, namely Sequential Attention (SA) (Yasuda et al., 2023), Sequential LASSO (SL) (Luo & Chen, 2014), Btach-wise Attenuation (LLY) (Liao et al., 2021), Group LASSO (GL) (Zhao et al., 2015), and Stochastic Gates (STG) (Yamada et al., 2020). [^3] These algorithms represent the top-performing methods reported in the literature. A comparison with other approaches (Atashgahi et al., 2022; Lemhadri et al., 2021; Sokar et al., 2022) demonstrates that SAND achieves better performance. For all methods except ours, training comprised a feature selection phase followed by a fitting phase wherein the neural network was retrained on the selected features. Although STG often eliminates the need for retraining thanks to its weight polarization, on some datasets (e.g. MICE Protein) the weights fail to polarize when $k=60$, even after an exhaustive search over the regularization parameter which indirectly governs the number of selected features, and so we resort to retraining in some cases. As for SAND, the fitting phase was omitted and the weights learned during the selection phase were directly utilized for inference. In other words, the gains corresponding to the non-selected features where set to zero while keeping all other weights of the model intact. Hence, from this point of view, our method offers two key benefits. Firstly, it demands fewer epochs ($33\%$ fewer epochs in our experiments). Secondly, it provides a streamlined pipeline where both selection and inference are handled by the same model.
@@
Figure 4: Test metrics for different $k$’s. Accuracy $(\uparrow)$ for all except MAE $(\downarrow)$ for CA Housing.
To provide additional insights, we varied the number of selected features $k$ under consistent settings and evaluated performance on the test set. Results are shown in Figure 4. Notably, SAND demonstrates its strength in feature selection, showcasing results that outperform or are comparable to other methods across different feature counts. This advantage is particularly significant given the fact that the best method is changing from dataset to dataset. Thus, there is a need for algorithms that deliver value beyond marginal accuracy improvements, prioritizing enhancements in computational demand and simplicity—qualities that our method exemplifies. It is important to note that all SAND experiments were conducted using fixed values of $\sigma=1.5$ and $\alpha=2$, without any fine-tuning. We omitted STG from Figure 4 due to the high computational cost of performing an exhaustive search over the regularization parameter $\lambda$ for each $k$. Moreover, on the MICE Protein dataset (77 features), STG fails to identify a clear feature subset once $k$ exceeds 37. For $k\leq 37$, STG produces a polarized solution—exactly $k$ features attain importance weight at almost 1, while the remainder are almost zero. However, for $k>37$, the selection process stalls: irrespective of $\lambda$, only the same 37 features remain at weight 1, and the other 40 features converge to an almost uniform, nonzero weight. Adjusting $\lambda$ cannot induce the selection of additional features; it only uniformly increases the weights of these 40 unselected features above zero.
@@
As discussed in Remark 2.2, the parameter $\sigma$ influences the rate of gain polarization. To demonstrate this, we trained SAND model on MICE dataset for 2000 epochs, using $\sigma$ values of 1.0 and 2.0, while keeping other settings fixed. We recorded the gains every 10 epochs. Figure 5 presents the sorted gains for selected epochs.We observe that while the gains tend to cluster around $1$ and 0 in both plots, this clustering occurs at a higher rate for a larger $\sigma$. Moreover, to assess SAND’s sensitivity to $\sigma$, we replicated the initial experiment while varying $\sigma\in\{1.0,1.5,2.0,2.5,3.0\}$. Results are presented in Table 1. As evident in the table, our approach demonstrates high robustness to the selection of $\sigma$, which underscores a positive aspect of the proposed method.
@@ 4 Summary and Future Works
  • On par with the state-of-the-art performance: Through extensive experiments, we showed that the proposed method has effectively state-of-the-art performance.
@@
Acknowledgements
The present work was developed within the AGRARSENSE Project that has received Chips JU funding (Grant Agreement No. 101095835). It has received funding from the Swiss State Secretariat for Education, Research and lnnovation (SERI) and is co-funded by the Innosuisse – Swiss Innovation Agency. More information: info@agrarsense.eu.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experimental set-up
Table 3: Dataset characteristics, experiment parameters, and all-features performance metrics. [^5] The metric is classification accuracy $(\uparrow)$ for all except MAE $(\downarrow)$ for CA Housing. Our study utilizes the entire MNIST and Fashion datasets, unlike related works. Additionally, the Activity dataset sourced from (Lemhadri et al., 2021)’s Google Drive and (Yasuda et al., 2023)’s repository contains 10,299 samples, as opposed to the 5,744 samples reported in the referenced papers.
@@
Moreover, the experiments were executed on a machine equipped with an NVIDIA GeForce RTX 4090 GPU with 24GB of RAM, paired with an AMD Ryzen 9 5900X 12-Core Processor featuring 24 threads. The code to reproduce our experiments is available at https://github.com/csem/SAND.
@@ Appendix B Experimental results
Table 4: Test metrics over 10 trials (mean $\pm$ standard deviation): Accuracy $(\uparrow)$ for all except MAE $(\downarrow)$ for CA Housing.
Dataset SA LLY GL SL STG SAND
Mice Protein $0.986\pm 0.005$ $0.983\pm 0.011$ $0.988\pm 0.007$ $0.987\pm 0.007$ $0.988\pm 0.007$ $\mathbf{0.988}\pm 0.006$
MNIST $0.961\pm 0.002$ $0.949\pm 0.003$ $0.930\pm 0.006$ $\mathbf{0.963}\pm 0.002$ $0.962\pm 0.002$ $0.962\pm 0.002$
MNIST-Fashion $0.836\pm 0.003$ $0.830\pm 0.004$ $0.828\pm 0.004$ $0.829\pm 0.003$ $\mathbf{0.856}\pm 0.002$ $0.836\pm 0.005$
ISOLET $0.927\pm 0.004$ $0.908\pm 0.011$ $0.924\pm 0.003$ $0.924\pm 0.007$ $\mathbf{0.936}\pm 0.005$ $0.926\pm 0.005$
COIL-20 $0.996\pm 0.005$ $0.998\pm 0.003$ $0.998\pm 0.003$ $0.996\pm 0.003$ $0.997\pm 0.004$ $\mathbf{0.998}\pm 0.003$
Activity $0.920\pm 0.009$ $0.894\pm 0.050$ $\mathbf{0.930}\pm 0.004$ $0.915\pm 0.005$ $0.922\pm 0.006$ $0.923\pm 0.004$
Madelon $0.786\pm 0.008$ $\mathbf{0.862}\pm 0.027$ $0.612\pm 0.010$ $0.603\pm 0.011$ $0.793\pm 0.03$ $0.732\pm 0.021$
HAR70 $0.901\pm 0.001$ $\mathbf{0.910}\pm 0.001$ $0.910\pm 0.003$ $0.908\pm 0.004$ $0.897\pm 0.002$ $0.897\pm 0.005$
CA Housing $0.639\pm 0.063$ $0.636\pm 0.063$ $0.590\pm 0.097$ $0.589\pm 0.097$ $0.577\pm 0.119$ $\mathbf{0.521}\pm 0.027$
@@
Table 5: Selection consistency on HAR70: feature selection frequency of the useful features (05)(0-5) and the misselected ones over 10 runs for $k=6$.
Feature SA SL SAND STG GL LLY
0 10 10 10 10 10 10
1 8 9 8 8 9 10
2 10 10 10 10 10 10
3 10 10 10 10 10 10
4 10 7 10 10 10 10
5 2 10 10 10 10 10
Misselected 10 4 2 2 1 0